Find the LCM of 40, 50, and 60. (show your work)
Thanks Precious

Answer by fractalier(6550)

Let us do a prime factorization of each first, so we have
40 = 2^3 * 5
50 = 2 * 5^2
60 = 2^2 * 3 * 5
Now we look for the greatest power of each of the prime numbers present in any of the factorizations...so we have
LCM = 2^3 * 3 * 5^2 = 600